A mother is 28 years older than her son,the mother's age is 10 years more than twice her son's age.what are their ages.
Answer by checkley79(3341)

Mother=s+26
Mother=2s+10
s+26=2s+10
s-2s=10-26
-s=-16
s=16 years old is the son's age.
16+26=42 years old is the mother's age.
Proof:
2*16+10=42
32+10=42
42=42